April Stigum will receive a 6 year annuity of $2,000 per year, beginning 7 years from today. In other words, the first payment will be made at the end of year 7. Assuming a required rate of return of 10%, calculate the present value today of her annuity.
Answer #1

Present value at year 6 = Annuity * [1 - 1 / (1 + r)n] / r

Present value at year 6 = 2000 * [1 - 1 / (1 + 0.1)6] / 0.1

Present value at year 6 = 2000 * [1 - 0.564474] / 0.1

Present value at year 6 = 2000 * 4.355261

Present value at year 6 = $8,710.5214

Present value today = Future value / (1 + r)n

Present value today = 8,710.5214 / (1 + 0.1)6

Present value today = 8,710.5214 / 1.771561

Present value today = $4,916.86